OpenStudy (anonymous):

You roll 2 dice. What is the probability of getting a 3 or 6 on the second die, given that you rolled a 1 on the first die?

OpenStudy (anonymous):

what is the formula that you are using? I am so confused on this!

OpenStudy (amistre64):

1 2 3 4 5 6 given -> 1 x x 2 3 4 5 6 since we are restricting ourselves to a "1"; there are only 6 options to choose from in that category; and a 3 or 6 are 2 of them sooo... 2/6 = 1/3

OpenStudy (amistre64):

the fancy stuff might be:\[P("3\ or\ 6"|1)\frac{P("3\ or\ 6"\ and\ 1)}{P(1)}\]

OpenStudy (amistre64):

i think my chart is more intuitive tho

OpenStudy (anonymous):

I am still a bit confused. If I have 12 chances on the first roll to get a 1, then I only have 2 chances to get a 3 or a 6, how does that really equate to 1/3. I'm not seeing it.

OpenStudy (amistre64):

12 chances? regardless, it says, given that you roll a 1, so that part is taken care of; now we focus within the set of (1,x), which there are only 6 ways total

OpenStudy (anonymous):

so I don't have to worry about the probability of rolling a 1? It's already figured out, so I just focus on the second part? I only see 2 possibilities for the second roll, either a 3 or a 6.

OpenStudy (amistre64):

the P(1) = 1/6 the P(3 or 6) = 2/6 the P("3 or 6" and 1) = 2/6 * 1/6 = 2/36 the P("3 or 6" | 1) = P("3 or 6" and 1) / P(1) 2/36 2/6 ---- = ---- = 2/6 = 1/3 1/ 6 1

OpenStudy (amistre64):

there are 36 total outcomes for this scenario: